# Synthesizing data products, mathematical models, and observational measurements for lake temperature forecasting

### v1

@@ Abstract
We present a novel forecasting framework for lake water temperature profiles, crucial for managing lake ecosystems and drinking water resources. The General Lake Model (GLM), a one-dimensional process-based model, has been widely used for this purpose, but, similar to many process-based simulation models, it: requires a large number of input variables, many of which are stochastic; presents challenges for uncertainty quantification (UQ); and can exhibit model bias. To address these issues, we propose a Gaussian process (GP) surrogate-based forecasting approach that efficiently handles large, high-dimensional data and accounts for input-dependent variability and systematic GLM bias. We validate the proposed approach and compare it with other forecasting methods, including a climatological model and raw GLM simulations. Our results demonstrate that our bias-corrected GP surrogate (GPBC) can outperform competing approaches in terms of forecast accuracy and UQ up to two weeks into the future.
@@ 1 Introduction
Process-based models are critical for understanding, managing, and forecasting environmental phenomena and are widely used in ecological research. The use of process-based models for one application in particular, the forecasting of lake water temperature profiles, is crucial for policies and decisions that impact the health of aquatic ecosystems (Carey et al., 2022; Cuddington et al., 2013). A lake temperature profile refers to the variation in water temperature at different depths within a lake. In aquatic ecosystems, water quality and the health of aquatic organisms is strongly associated with water temperature (Carey et al., 2022), which influences various biological processes, such as the growth and distribution of aquatic flora and fauna (Wetzel, 2001). Warmer water temperatures can lead to accelerated growth of harmful phytoplankton, which can produce toxins that threaten aquatic and human health (Carey et al., 2012). Lake temperature forecasts that extend at least 30 days into the future, include uncertainty quantification (UQ), and are updated frequently are crucial to adequately managing drinking water resources, mitigating water quality degradation, and planning for the health of lake ecosystems (Thomas et al., 2023; Carey et al., 2022). In this study, we focus on forecasting lake temperatures at ten depths at Falling Creek Reservoir (FCR), a small drinking water reservoir located in Vinton, Virginia, USA (Figure 1 left).
@@
There are several approaches for forecasting lake temperatures (Lofton et al., 2023), the most simple of which are climatological forecasts. These are based on historical data, capturing broader trends, and are accurate at long (15+ day) horizons (Thomas et al., 2023). Such forecasts for FCR can be constructed using data collected by in situ water temperature sensors (denoted by the red dot in Figure 1 left). By definition, climatological forecasts do not vary substantively year upon year, predictive limits do not vary with the forecast horizon, and they lack sharpness (the tendency of a model to predict outside of the mean Gneiting, Balabdaoui and Raftery, 2007), reducing their value for near-term water quality management.
@@
Improved lake temperatures forecasts can be obtained by driving a process-based lake ecosystem model with an ensemble of weather forecasts (Cho et al., 2020; Thomas et al., 2020, 2023). Lake ecosystem models are designed to simulate heat transfer within aquatic ecosystems (Mooij et al., 2010) and can encompass a wide range of processes, including hydrodynamics (Hipsey et al., 2019), water quality (Hipsey et al., 2012; Cole and Wells, 2021), nutrient cycling (Janse, Aldenberg and Kramer, 1992), and ecological interactions (Leon et al., 2006). One such model, the one-dimensional (1D) General Lake Model (GLM; Hipsey et al., 2019) [^1] The authors apologize for this overloaded acronym; throughout “GLM” refers to the General Lake Model, not a generalized linear model. is a general-purpose lake ecosystem model that can simulate thermal dynamics in a vertical column of a water body. GLM takes into account the physical characteristics of the lake, such as its depth, surface area, and geometry, water inputs from surrounding areas, and external factors (e.g., meteorology) to simulate water movement, mixing, and circulation patterns within a column (Hipsey et al., 2019). GLM has been used to forecast long-term lake ecosystem responses to climate change by driving it with climate projections (Christianson, Johnson and Hooten, 2020; Gal et al., 2020; Fenocchi et al., 2018). It has also been used to study the effects of climate change on long-term historical lake dynamics (Ladwig et al., 2018; Shikhani et al., 2022) as well as to research general historical lake hydrodynamics (Huang et al., 2017; Bueche, Hamilton and Vetter, 2017) and water quality indicators (Ladwig et al., 2021; Ward et al., 2020).
@@
However, GLM has characteristics that make it challenging to use for near-term (1-30 days into the future) forecasts of lake temperatures. To obtain lake temperature forecasts, GLM must be driven with forecasted weather products. Furthermore, since GLM is a deterministic model, UQ associated with forecasts can only be derived if it is driven with an ensemble of weather forecasts or parameter values, although we do not consider the latter here. Importantly, like all process-based models, GLM is a simplified version of reality that does not perfectly represent processes that govern energy transfer within a water column, let alone in an entire (three-dimensional) lake. Any systematic biases, or other errors that result from such oversimplifications, need to be accounted for.
Observations provide a critical foundation for assessing the accuracy and reliability of GLM. At FCR, water temperature sensors installed at 10 depths (surface to 9m depth by 1m increments) have collected a rich and continuous observational data record starting in July 2018 (red dot in Figure 1, left). One can get a sense of GLM performance for the FCR study region by comparing GLM output to sensor data. Figure 1 (right) shows an example of GLM simulations (yellow) and sensor observations (black) at 6m depth (a depth near the thermocline where model performance is often lowest relative to other depths) as well as a surrogate model for GLM in blue (discussed momentarily). These GLM simulations exhibit a pronounced warm bias between April and October with respect to observations. Although the bias appears obvious via a visual comparison of raw GLM output and observations, it is not possible to visually separate GLM bias from observational noise. To do so, an additional model for the bias itself is necessary, and we expand upon this idea in the following sections.
@@
Figure 1: Left: Falling Creek Reservoir (FCR) in Vinton, Virginia, USA (37.30N, 79.84W). The location of sensors, installed in 2018, are denoted by the red dot on the map and are located at 10 depths (surface to 9m at 1-m increments). Lake depth (m) is shown by the blue color scale. Right: Example GLM simulations (yellow) and observations (black) for 2021 at a lake depth of 6m. Means and 90% prediction intervals of a GLM surrogate are shown in blue; a bias-corrected version is shown in pink.
@@
We seek accurate 1-30 day ahead ensemble forecasts of lake temperatures at 10 depths at FCR with well-calibrated UQ. These forecasts are made by driving GLM with a 31-member ensemble of weather forecasts generated by the National Oceanic and Atmospheric Administration Global Ensemble Forecast System (NOAA-GEFS; Hamill et al., 2022, see https://www.ncei.noaa.gov/products/weather-climate-models/global-ensemble-forecast). Construction of ensemble lake temperature forecasts is facilitated with automated code (available at https://github.com/maikeh7/Surrogate_Assisted_GLM) and is inspired by Thomas et al. (2020). Figure 2 diagrams how NOAA ensemble forecasts and data from FCR are used to construct lake temperature forecasts with GLM.
Figure 2: This schematic shows how NOAA forecasts and data from FCR are utilized to generate 1–30 day ahead forecasts at 10 depths for the time period of interest (2020-10-03 and 2023-06-10). NOAA 1–30 day ahead forecasts consist of 31 ensemble members of 4 variables: precipitation, relative humidity, surface temperature, and solar radiation. Temperature sensors at FCR (red dot) collect data at 10 depths (surface to 9m depth at 1-m intervals). To generate lake temperature forecasts for FCR for the 900 reference dates in 2020-10-03 and 2023-06-10, GLM requires weather forecast inputs as well as site-dependent features such as lake size, area, sediment type, and general morphology. Given the weather inputs and site-dependent parameters from FCR, GLM can generate 31, 1 to 30 day-ahead forecasts of lake temperatures at 10 depths. Each of the 31 resulting forecasts can be visualized as a 2D surfaces of lake temperature over horizon ($h$) and depth ($d$). Observed data from FCR can be visualized as 2D scatter plots of lake temperature over $h$ and $d$. It is important to note that the amount of observed data is substantially lower than those generated by GLM.
@@
Examples of NOAA-driven GLM forecasts (hereafter “NOAA-GLM”) at FCR for 1–30 days in the future are shown in Figure 3. At lower horizons, the forecasts of individual ensemble members are similar, but as horizon increases, forecasts diverge. There is considerable variability in within reference dates: forecasts starting on 2020-10-30 (fall, Figure 3A) are less variable compared to forecasts originating on 2021-01-01 (winter, Figure 3 B). Variability also differs according to lake depth, where shallow depths exhibit more variation than deeper ones. While simple, this approach is useful: forecasts have long (30-day) horizons, and uncertainty can be derived from variability among ensemble members. However, it does not account for GLM bias, and uncertainty is not modeled directly, making UQ less reliable.
@@
Figure 3: NOAA-GLM lake temperature forecasts (C) for FCR. Forecasts start on reference dates A) 2020-10-30 and B) 2021-01-01. Colors of lines denote trajectories of individual ensemble members. Panel numbers denote lake depth (m).
@@
A more robust alternative to generating lake temperature forecasts involves building a probabilistic representation of NOAA-GLM, which could subsequently improve prediction accuracy and UQ, while also denoising the 31 ensemble forecast trajectories. This approach could accommodate functional forms of bias, allowing for a flexible bias-correction. Modeling bias is also advantageous because it allows the bias to be separated from observational noise (Liu et al., 2009); as we saw in Figure 1 (right), this is not possible using raw model output and observations. Observe in Figure 1 how one type of statistical model (blue lines) adequately smooths GLM simulations while providing appropriate UQ. While the blue lines represent a statistical model for only a single GLM realization, we propose to extend this idea to modeling GLM ensemble forecast trajectories. Figure 1 (right) also shows the mean and 90% prediction intervals resulting from a bias-corrected model (pink lines). In contrast to the biased model (blue) fitted to GLM simulations, the mean of the bias-corrected model provides a very good approximation to observations, and UQ is reliable.
@@
One way to statistically model NOAA-GLM forecasts is through the use of a computer model surrogate. A surrogate emulates the behavior of a computer model such as GLM but does so without the model’s physical equations in its implementation (Santner et al., 2018). Gaussian process (GP) surrogates are the canonical choice as they have many desirable properties, including flexibility, smoothness, predictive accuracy, and reliable UQ (Gramacy, 2020; Williams and Rasmussen, 2006). Model bias can also be modeled with a GP (Kennedy and O’Hagan, 2001; Liu et al., 2009), allowing for robust and versatile bias-correction. One major challenge to using a GP surrogate is the decomposition bottleneck of multivariate normal (MVN) covariance matrices. In GP fitting/prediction, these matrices are sized commensurately with the square of the training data of size $n$, incurring a decomposition cost that is cubic in $n$ flops. In our application, $n$ is very large. Each of the 31 NOAA ensemble members results in 31 NOAA-GLM forecasts for 30 horizons and 10 lake depths, producing a total dataset size of: 900 reference dates $\times$ 30 horizons $\times$ 10 depths $\times$ 31 ensemble members = 8,370,000, a formidable data size for any model-fitting enterprise, and especially so for GP surrogates.
GP surrogates have been used in hydrological (Razavi, Tolson and Burn, 2012) and environmental applications (Fer et al., 2018), model calibration (Fer et al., 2018), sensitivity (Oakley and O’Hagan, 2004) and uncertainty analysis (Yang et al., 2018; Roy et al., 2018). However, bias-corrected, GP surrogates that can handle input-dependent variability and big data have never been implemented in the context of ecological forecasting. The main contribution of this paper is the development of a novel lake temperature forecasting workflow for water quality management, where NOAA-GLM forecasts are represented by a bias-corrected GP surrogate. Importantly, our approach, which can generate forecasts at 10 depths and 30 forecast horizons, is efficient enough to allow for iterative forecasting, where forecasts, models and datasets are updated daily.
@@
This paper is organized as follows: we begin with a review of GP surrogates for computer models and follow with a detailed description of the forecasting workflow and the statistical techniques associated with it. We validate our proposed framework and show how forecasts from our framework compare to one in which bias is not corrected, raw GLM output, and a climatological forecast. We then discuss results and provide concluding remarks.
@@ 2 Background and proof-of-concept
Our forecasting framework operates on the frontier of Gaussian process (GP) surrogate modeling (e.g., Gramacy, 2020) for synthesizing computer model output, observations, and correcting bias. A basic GP setup is introduced here, building upon Figure 1. That earlier analysis comprises a proof-of-concept, suggesting that forecasts utilizing GLM and lake temperatures measured by sensors would benefit from meta-modeling and bias correction. Ultimately, we will need a higher-powered strategy in order to cope with a large corpus of NOAA-GLM runs, which is the subject of Section 3.
@@ 2.1 Gaussian process surrogate modeling
GPs are popular as surrogates for computer model simulation experiments (e.g., Santner et al., 2018), and more widely as a nonlinear and nonparametric regression tool (e.g., Williams and Rasmussen, 2006), because they furnish accurate predictors with appropriate coverage when used to model smooth and stationary (i.e., physical) phenomena. Here we utilize GPs in both contexts: for computer model simulations and on observational data. As models for computer experiments, GP surrogates can be crucial to enabling subsequent tasks such as optimization, calibration, and sensitivity analysis (e.g., Gramacy, 2020). One advantage GPs enjoy over other nonlinear regression choices – such as splines (Friedman, 1991), polynomial regression (Jones, 2001), and neural networks (Farfán and Cea, 2021) – is well-calibrated, probabilistic uncertainty quantification (UQ), which is easy to filter to the “subsequent tasks” listed above. This UQ extends to input-dependent variability (e.g., Ankenman, Nelson and Staum, 2008; Binois, Gramacy and Ludkovski, 2018), such as is seen in GLM forecasts over increasing horizon [Figure 3] and is discussed in greater detail below in Section 3.
Consider a computer model $M$ represented as a function $f:\mathbb{R}^{p}\rightarrow\mathbb{R}$ that maps inputs to outputs, whose pairs over $n_{\_}M$ runs of a simulation campaign comprise data $D_{\_}{n_{\_}M}=(\mathbf{x}_{\_}i,y_{\_}i)$, $i=1\dots n_{\_}M$. Let $\mathbf{X}_{\_}{n_{\_}M}=(\mathbf{x}_{\_}1^{\top},\mathbf{x}_{\_}2^{\top},\dots,\mathbf{x}_{\_}{n_{\_}M}^{\top})\in\mathbb{R}^{n_{\_}M\times p}$, collect inputs, and $\mathbf{Y}^{M}_{\_}{n_{\_}M}=\left(y_{\_}1,y_{\_}2,\dots,y_{\_}{n_{\_}M}\right)^{\top}$ outputs. Utilizing a GP model, or prior, for $f$ amounts to specifying that outputs $\mathbf{Y}_{\_}{n_{\_}M}^{M}$ jointly follow a multivariate normal distribution (MVN). In a GP regression context it is common that the MVN be specified with a mean of zero, and a covariance $\Sigma$ that is determined by (inverse) distances between inputs $\mathbf{X}_{\_}{n_{\_}M}$:
@@
Our setup is largely indifferent to variations in choices of $q$, $k$, etc. In our empirical work we use the $q$ specified in Eq. (2), and the Matern kernel $k$ with smoothness level fixed at 3.5.
@@
$L(\boldsymbol{\theta};D_{\_}{n_{\_}M})\propto|\Sigma_{\_}{\boldsymbol{\theta}}|^{-1/2}\mathrm{exp}\left(-\frac{1}{2}(\mathbf{Y}^{M}_{\_}{n_{\_}M})^{\top}\Sigma_{\_}{\boldsymbol{\theta}}^{-1}\mathbf{Y}^{M}_{\_}{n_{\_}M}\right)$ (3)
which can be used to learn settings $\hat{\boldsymbol{\theta}}_{\_}{n_{\_}M}$; see, e.g., Gramacy (2020, Chapter 5). Observe that $\boldsymbol{\theta}$ is buried in $\Sigma_{\_}{\boldsymbol{\theta}}$, although we often drop the subscript to streamline the notation. It is common in the literature to presume that settings $\hat{\boldsymbol{\theta}}_{\_}{n_{\_}M}$ have already been estimated given data $D_{\_}{n_{\_}M}$ and focus analysis on other aspects of inference. More detail on the nuances of our own estimation procedures for an expanded hyperparameter $\boldsymbol{\theta}$ is provided in Section 3. However, it is worth pointing out that determinants and inverses, as required to evaluate the likelihood (3) require cubic in $n_{\_}M$ flops for dense matrices $\Sigma_{\_}{\boldsymbol{\theta}}$, a computational bottleneck for large $n_{\_}M$. Training data sizes in the low thousands is manageable; tens of thousands require minutes or hours, or are entirely infeasible.
One of the most important uses for GPs involves prediction. When modeling computer simulations, the predictive equations are what are referred to as the “surrogate”, because these $\hat{f}(\cdot)$ may be used in lieu of a new simulation $f(\cdot)$. In geostatistics these are sometimes called the kriging equations (Banerjee, Carlin and Gelfand, 2003). Suppose we wish to predict $\mathbf{Y}_{\_}{n_{\_}M}$ at a new $n^{\prime}\times d$ set of input configurations collected in the rows of $\mathcal{X}$. Deriving the distribution for $\mathbf{Y}(\mathcal{X})$ involves stacking the MVN for $\mathbf{Y}_{\_}{n_{\_}M}$ together with an analogous one for $\mathbf{Y}(\mathcal{X})$, i.e., where the latter follows Eq. (1) but where $\Sigma(\mathcal{X})$ is used instead. The joint, $(n_{\_}M+n^{\prime})$-dimensional MVN has a covariance structure with block diagonal components $\Sigma(\mathbf{X}_{\_}{n_{\_}M})$ and $\Sigma(\mathcal{X})$, and off-diagonal $\Sigma(\mathbf{X}_{\_}{n_{\_}M},\mathcal{X})$ and its transpose. The quantity $\Sigma(\mathbf{X}_{\_}{n_{\_}M},\mathcal{X})$ is derived from (inverse) distances between training $\mathbf{X}_{\_}{n_{\_}M}$ and testing $\mathcal{X}$ but involves no nugget ($g$) augmentation. Then, using MVN conditioning identities, we may derive the predictive/conditional distribution $\mathbf{Y}(\mathcal{X})\mid D_{\_}{n_{\_}M}\sim\mathcal{N}_{\_}{n^{\prime}}(\mu_{\_}{n_{\_}M}(\mathcal{X}),\Sigma_{\_}{n_{\_}M}(\mathcal{X}))$, where
@@
If, instead, one is interested in predictions for $f(\mathcal{X})=\mathbb{E}\{\mathbf{Y}(\mathcal{X})\}$ one would take $g=0$ in $k(\mathcal{X})+g\mathbb{I}_{\_}{n^{\prime}}$ in Eq. (4) to obtain confidence intervals (CIs) instead. More specifically, and since we will need to use it later, if $\boldsymbol{\mu}_{\_}{n_{\_}M}\equiv\mu_{\_}{n_{\_}M}(\mathcal{X})=\mathbb{E}\{\mathbf{Y}_{\_}{n_{\_}M}(\mathcal{X})\}$, then

### v2

# Synthesizing data products, mathematical models, and observations for lake temperature forecasting

@@ Abstract
We present a novel forecasting framework for lake water temperature, which is crucial for managing lake ecosystems and drinking water resources. The General Lake Model (GLM) has been previously used for this purpose, but, similar to many process-based simulation models, it: requires a large number of inputs, many of which are stochastic; presents challenges for uncertainty quantification (UQ); and can exhibit model bias. To address these issues, we propose a Gaussian process (GP) surrogate-based forecasting approach that efficiently handles large, high-dimensional data and accounts for input-dependent variability and systematic GLM bias. We validate the proposed approach and compare it with other forecasting methods, including a climatological model and raw GLM simulations. Our results demonstrate that our bias-corrected GP surrogate (GPBC) can outperform competing approaches in terms of forecast accuracy and UQ up to two weeks into the future.
@@ 1 Introduction
Process-based models are critical for understanding, managing, and forecasting environmental phenomena and are widely used in ecological research (Cuddington et al., 2013). They are crucial for forecasting lake water temperature profiles (e.g., at varying depth) that influence decisions that impact the health of aquatic ecosystems (Carey et al., 2022). Water quality and the health of aquatic organisms are strongly associated with water temperature (Carey et al., 2022), which influences the growth and distribution of aquatic flora and fauna (Wetzel, 2001). Warmer water can accelerate the growth of harmful phytoplankton, threatening aquatic and human health (Carey et al., 2012). Lake temperature forecasts that extend at least 30 days into the future, include uncertainty quantification (UQ), and are updated frequently are crucial to adequately managing drinking water resources, mitigating water quality degradation, and planning for the health of lake ecosystems (Thomas et al., 2023; Carey et al., 2022). Here, we focus on forecasting lake temperatures at ten depths at Falling Creek Reservoir (FCR), a small reservoir in Vinton, Virginia, USA (Figure 1 left).
@@
There are several approaches to forecasting lake temperatures (Lofton et al., 2023), the simplest of which are climatological. These are based on historical data, capturing broader trends, and are accurate at long (15+ day) horizons (Thomas et al., 2023). Such forecasts for FCR can be constructed using data collected by in situ water temperature sensors (denoted by the red dot in Figure 1 left). Climatological forecasts do not vary substantively year upon year, their predictive limits do not vary with the forecast horizon, and they lack sharpness (Gneiting, Balabdaoui and Raftery, 2007), reducing their value for near-term water quality management.
@@
Lake ecosystem models are designed to simulate heat transfer within aquatic ecosystems (Mooij et al., 2010) and can encompass a wide range of processes, including hydrodynamics (Hipsey et al., 2019), water quality (Hipsey et al., 2012; Cole and Wells, 2021), nutrient cycling (Janse, Aldenberg and Kramer, 1992), and ecological interactions (Leon et al., 2006). One such model, the one-dimensional (1D) General Lake Model (GLM; Hipsey et al., 2019), [^1] Apologies for the acronym; here “GLM” refers to the General Lake Model, not a generalized linear model. can simulate thermal dynamics in a vertical column of a waterbody. GLM takes into account the physical characteristics of the lake, such as its depth, surface area, and geometry, water inputs from surrounding areas, and external factors (e.g., meteorology) to simulate water movement, mixing, and circulation patterns within the water column (Hipsey et al., 2019). GLM has been used to predict long-term lake ecosystem responses to climate change with climate projection driver data (Christianson, Johnson and Hooten, 2020; Gal et al., 2020; Fenocchi et al., 2018). It has also been used to study the effects of climate change on historical lake dynamics (Ladwig et al., 2018; Shikhani et al., 2022) as well as to study lake hydrodynamics (Huang et al., 2017; Bueche, Hamilton and Vetter, 2017) and water quality indicators (Ladwig et al., 2021; Ward et al., 2020).
@@
However, the GLM has characteristics that make it challenging to use for near-term (1–30 days) forecasts of lake temperatures. Since GLM is a deterministic model, UQ associated with forecasts can only be derived if it is driven with an ensemble of model inputs, such as weather forecasts or model parameters. Importantly, like all process-based models, GLM is a simplified version of reality that does not perfectly represent processes that govern energy transfer within a water column, let alone in an entire (three-dimensional) lake. Any systematic biases, or other errors that result from such oversimplifications, need to be accounted for.
@@
Figure 1: Left: Falling Creek Reservoir (FCR; 37.30N, 79.84W). Red dot denotes sensor location. Lake depth (m) is shown by the graduated color scale. Right: Example GLM simulations (thick yellow) and observations (thick black) for 2021 at 6m; means and 90% PIs of a GLM surrogate (thinner blue); a bias-corrected (thinner pink).
Observations provide a critical foundation for assessing the accuracy of GLM. At FCR, water temperature sensors installed at 10 depths (surface to 9m depth by 1m increments) provide rich and continuous observational data record starting in July 2018 (red dot in Figure 1 left). One can get a sense of GLM performance for the FCR study region by comparing GLM output to sensor data. Figure 1 (right) shows an example of GLM simulations (thick yellow line) and sensor observations (thick black line) at 6m depth as well as a surrogate model for GLM in thinner blue lines (discussed momentarily). These GLM simulations exhibit a pronounced warm bias between April and October with respect to observations.
@@
We seek accurate 1-30 day ahead ensemble forecasts of lake temperatures at 10 depths at FCR with well-calibrated UQ. These forecasts are made by driving GLM with a 31-member ensemble of weather forecasts generated by the National Oceanic and Atmospheric Administration Global Ensemble Forecast System (NOAA-GEFS; Hamill et al., 2022, see https://www.ncei.noaa.gov/products/weather-climate-models/global-ensemble-forecast). Construction of ensemble lake temperature forecasts is facilitated with automated, publicly available code (https://github.com/maikeh7/Surrogate_Assisted_GLM) and is inspired by Thomas et al. (2020). Figure 2 diagrams how NOAA ensemble forecasts and data from FCR are used to construct lake temperature forecasts with GLM.
Figure 2: Schematic relating NOAA forecasts to GLM simulations to FCR sensor data sensor data towards deriving 1–30 day ahead forecasts at 10 depths from 2020-10-03 to 2023-06-10.
@@
Examples of NOAA-driven GLM forecasts (hereafter “NOAA-GLM”) at FCR for 1–30 days in the future are shown in Figure 3. At lower horizons, the forecasts of individual ensemble members are similar, but as horizon increases, forecasts diverge. There is considerable variability within reference dates: forecasts starting on 2020-10-30 (fall, Figure 3A) are more variable compared to forecasts originating on 2021-01-01 (winter, Figure 3B). Variability also differs according to lake depth, where shallow depths exhibit more variation than deeper ones. While simple, this approach is useful: forecasts have long (30-day) horizons, and uncertainty can be derived from variability among ensemble members. However, it does not account for GLM bias, and uncertainty is not modeled directly, making UQ less reliable.
@@
Figure 3: NOAA-GLM lake temperature forecasts (C) for FCR starting on reference dates A) 2020-10-30 and B) 2021-01-01. Colors denote trajectories of individual ensemble members. Panel numbers denote depth (m).
@@
A more robust alternative to generating lake temperature forecasts involves building a probabilistic representation of NOAA-GLM, which could subsequently improve prediction accuracy and UQ, while also denoising the 31 ensemble forecast trajectories. This approach could accommodate functional forms of bias and correction. Modeling bias is also advantageous because it allows the bias to be separated from observational noise (Liu et al., 2009). Observe in Figure 1 how one type of statistical model (blue lines) adequately smooths GLM simulations while providing appropriate UQ. We propose to extend this idea from single to ensemble GLM forecast trajectories. Figure 1 (right) also shows the mean and 90% prediction intervals resulting from a bias-corrected model (pink lines). In contrast to the biased model (thin blue lines) fitted to GLM simulations, the mean of the bias-corrected model provides a very good approximation to observations, and UQ is reliable.
@@
One way to statistically model NOAA-GLM forecasts is through the use of a computer model surrogate. A surrogate emulates the behavior of a computer model such as GLM but does so without the model’s physical equations (Santner et al., 2018). Gaussian process (GP) surrogates are the canonical choice as they have many desirable properties, including flexibility, accuracy, and reliable UQ (Gramacy, 2020; Williams and Rasmussen, 2006). Bias can also be modeled with a GP (Kennedy and O’Hagan, 2001; Liu et al., 2009). One major challenge to using a GP surrogate is the decomposition bottleneck of multivariate normal (MVN) covariance matrices. In GP fitting/prediction, these matrices are sized commensurately with the square of the training data of size $n$, incurring a decomposition cost that is cubic in $n$ flops. In our application, $n$ is very large. Each of the 31 NOAA ensemble members results in 31 NOAA-GLM forecasts for 30 horizons and 10 lake depths, producing a total dataset size of: 900 reference dates $\times$ 30 horizons $\times$ 10 depths $\times$ 31 ensemble members $\rightarrow n=8{,}370{,}000$, a formidable data size for any model-fitting enterprise, and especially so for GP surrogates.
@@
GP surrogates have been used in hydrological (Razavi, Tolson and Burn, 2012) and environmental applications (Fer et al., 2018), model calibration (Fer et al., 2018), sensitivity (Oakley and O’Hagan, 2004) and uncertainty analysis (Yang et al., 2018; Roy et al., 2018). However, bias-corrected, GP surrogates that can handle input-dependent variability and big data have never been implemented in the context of ecological forecasting. We aim to generate bias-corrected forecasts (10 depths and 30 horizons) via NOAA-GLM ensembles fast enough so updates for all quantities may be furnished daily. The remainder of the paper is organized as follows: we begin with a review of GP surrogates for computer models and follow with a detailed description of the forecasting workflow and the statistical techniques associated with it. We validate our proposed framework and show how forecasts from our framework compare to one in which bias is not corrected, raw GLM output, and a climatological forecast. We then discuss the results and provide concluding remarks.
@@ 2 Background and proof-of-concept
Our forecasting framework operates on the frontier of Gaussian process (GP) surrogate modeling (e.g., Gramacy, 2020) for synthesizing computer model output, observations, and correcting bias. A basic GP setup is introduced here, building upon Figure 1. This preliminary analysis suggests that forecasts utilizing GLM and lake temperatures measured by sensors would benefit from meta-modeling and bias correction. Ultimately, we will need a higher-powered strategy in order to cope with a large corpus of NOAA-GLM runs, which is the subject of Section 3.
@@ 2.1 Gaussian process surrogate modeling
GPs are popular as surrogates for computer model simulations (e.g., Santner et al., 2018), and more widely for nonlinear regression (e.g., Williams and Rasmussen, 2006), because they furnish accurate predictors with appropriate coverage when used to model smooth and stationary (i.e., physical) phenomena. Here, we utilize GPs in both contexts: for computer model simulations and for observational data.
Consider a computer model $M$ represented as a function $f:\mathbb{R}^{p}\rightarrow\mathbb{R}$ that maps inputs to outputs, whose pairs over $n_{\_}M$ simulations comprise data $D_{\_}{n_{\_}M}=(\mathbf{x}_{\_}i,y_{\_}i)$, $i=1\dots n_{\_}M$. Let $\mathbf{X}_{\_}{n_{\_}M}=(\mathbf{x}_{\_}1^{\top},\mathbf{x}_{\_}2^{\top},\dots,\mathbf{x}_{\_}{n_{\_}M}^{\top})\in\mathbb{R}^{n_{\_}M\times p}$, collect inputs, and $\mathbf{Y}^{M}_{\_}{n_{\_}M}=\left(y_{\_}1,y_{\_}2,\dots,y_{\_}{n_{\_}M}\right)^{\top}$ outputs. Utilizing a GP model, or prior, for $f$ amounts to specifying that $\mathbf{Y}_{\_}{n_{\_}M}^{M}$ follow a multivariate normal distribution (MVN). It is common that the MVN be specified with a mean of zero, and a covariance $\Sigma$ that is determined by (inverse) distances between inputs $\mathbf{X}_{\_}{n_{\_}M}$:
@@
Our setup is largely indifferent to variations in choices of $q$, $k$, etc. In our empirical work we use the $q$ specified in Eq. (2), and the Matérn kernel $k$ with smoothness level fixed at 3.5.
@@
$L(\boldsymbol{\theta};D_{\_}{n_{\_}M})\propto|\Sigma_{\_}{\boldsymbol{\theta}}|^{-1/2}\mathrm{exp}\left(-\frac{1}{2}(\mathbf{Y}^{M}_{\_}{n_{\_}M})^{\top}\Sigma_{\_}{\boldsymbol{\theta}}^{-1}\mathbf{Y}^{M}_{\_}{n_{\_}M}\right),$ (3)
which can be used to learn settings $\hat{\boldsymbol{\theta}}_{\_}{n_{\_}M}$; see, e.g., Gramacy (2020, Chapter 5). Observe that $\boldsymbol{\theta}$ is buried in $\Sigma_{\_}{\boldsymbol{\theta}}$, although we often drop the subscript to streamline notation. We presume $\hat{\boldsymbol{\theta}}_{\_}{n_{\_}M}$ has been pre-estimated given data $D_{\_}{n_{\_}M}$ to focus on other aspects of inference. More detail on our own procedures for an expanded $\boldsymbol{\theta}$ is provided in Section 3. Note that determinants and inverses in Eq. (3) require cubic in $n_{\_}M$ flops for dense matrices $\Sigma_{\_}{\boldsymbol{\theta}}$.
Suppose we wish to predict outputs at a new $n^{\prime}\times d$ set of input configurations collected in the rows of $\mathcal{X}$. For computer simulations, predictive equations comprise the “surrogate”, because these $\hat{f}(\cdot)$ may be used in lieu of a new simulation $f(\cdot)$. In geostatistics these are known as the kriging equations (Banerjee, Carlin and Gelfand, 2003). Deriving the distribution for $\mathbf{Y}(\mathcal{X})$ involves stacking the MVN for $\mathbf{Y}_{\_}{n_{\_}M}$ together with an analogous one for $Y(\mathcal{X})$, i.e., where the latter follows Eq. (1) but where $\Sigma(\mathcal{X})$ is used instead. The joint, $(n_{\_}M+n^{\prime})$-dimensional MVN has a covariance structure with block diagonal components $\Sigma(\mathbf{X}_{\_}{n_{\_}M})$ and $\Sigma(\mathcal{X})$, and off-diagonal $\Sigma(\mathbf{X}_{\_}{n_{\_}M},\mathcal{X})$ and its transpose. The quantity $\Sigma(\mathbf{X}_{\_}{n_{\_}M},\mathcal{X})$ is derived from distances between training $\mathbf{X}_{\_}{n_{\_}M}$ and testing $\mathcal{X}$ but involves no nugget ($g$) augmentation. Then, MVN conditioning provides $Y(\mathcal{X})\mid D_{\_}{n_{\_}M}\sim\mathcal{N}_{\_}{n^{\prime}}(\mu_{\_}{n_{\_}M}(\mathcal{X}),\Sigma_{\_}{n_{\_}M}(\mathcal{X}))$, where
@@
If, instead, one is interested in predictions for $f(\mathcal{X})=\mathbb{E}\{Y(\mathcal{X})\}$ one would take $g=0$ in $k(\mathcal{X})+g\mathbb{I}_{\_}{n^{\prime}}$ in Eq. (4) to obtain confidence intervals (CIs) instead. More specifically, and since we will need to use it later, if $\boldsymbol{\mu}_{\_}{n_{\_}M}\equiv\mu_{\_}{n_{\_}M}(\mathcal{X})=\mathbb{E}\{Y(\mathcal{X})\mid D_{\_}{n_{\_}M}\}$, then